OpenStudy (abbycross167):

Can someone please help me with this question? |- 3 ½ | - | 12 ¼ |

OpenStudy (anonymous):

Put them as top heavy fractions first so...\[-\frac{ 7 }{ 2 } - \frac{ 49 }{ 4 }\] Then make them have the same denominator... \[-\frac{ 14 }{ 4 }-\frac{ 49 }{ 4}\] So it's...\[-\frac{ 63 }{ 4}\]

OpenStudy (abbycross167):

thats it?

OpenStudy (anonymous):

Yep it is.. :)

OpenStudy (anonymous):

Unless the lines are Modulus lines? or are they brackets?